OpenStudy (lillian_a):

Find the equation of the line tangent to the graph of y= x^3-5x^2-2x when x = 2.

OpenStudy (bibby):

can you find f'(x)?

OpenStudy (lillian_a):

how?

OpenStudy (bibby):

take the derivative of the function

OpenStudy (lillian_a):

okay

OpenStudy (lillian_a):

after i find the y what do i do?

OpenStudy (bibby):

plug in x=2

OpenStudy (lillian_a):

i did and i found the y

OpenStudy (daniel.ohearn1):

After you find the corresponding y value you plug it in to the equation y-y1=m(x-x1) for y1. Then you can plug in 2 into the derivative represented as m, (the slope of the tangent line). Your answer will be in the form y=mx+b

OpenStudy (lillian_a):

do i plug x in to the original equation or the derivative to figure out the y?

OpenStudy (daniel.ohearn1):

The original equation to get y. You will then have a point. After you take the derivative you will have your m slope in the equation for a tangent line. Plug x=2 into it finally to simplify the slope.

OpenStudy (lillian_a):

Okay thank you

OpenStudy (daniel.ohearn1):

You're welcome!